Assuming that both numbers are in base 3, find the product of 21 x 21.
Answer by stanbon(57361)

Assuming that both numbers are in base 3, find the product of 21 x 21.
21 = 2*3+1 = 7
-------------------
7*7 = 49 = 3^3 + 2*3^2 + 3 + 1
---------------------
So the answer is base 3 is 1211
======================
Cheers,
Stan H.